# Two-Step Regularized HARX to Measure Volatility Spillovers in Multi-Dimensional Systems

### v3

@@ 1 Introduction
Third, Joint Impulse Response Functions (JIRFs) allow us to show how volatility shocks transmit and accumulate at multiple horizons. While the cross effects on our hybrid model were all near zero, the JIRFs allow us to show that in some cases, like equity shocks on crude oil futures realized volatility, we observed a sharp widening of confidence intervals. This demonstrates how volatility spillovers contribute to ’volatility of volatility’.
@@ 3.4 HAR Benchmark
3.5 Spillover Network Identification
The advantage of the hybrid HARX-ElasticNet framework is that the estimated coefficient matrix directly reveals the spillover network structure. The $\gamma$ coefficients indicate which cross-market transmission channels are economically meaningful (those that survive regularization) versus which are noise that gets shrunk to zero.
@@ 4.1 Forecast Performance
4.2 Spillover Network Structure
While HAR excels at point forecasting, it can’t identify cross-market volatility transmission. The HARX-ElasticNet reveals which spillover pathways are economically meaningful by shrinking noise to zero while preserving significant cross-market effects. Figure 3 presents the complete estimated coefficient matrix (18 features $\times$ 6 assets = 108 coefficients), where each element represents how lagged volatility in one market (source) at a specific frequency (daily, weekly, monthly) affects current volatility in another market (target).
@@ 4.2 Spillover Network Structure
Methodological Insight. The extreme sparsity of the hybrid model compared to standard HARX-ElasticNet highlights that applying regularization uniformly to all coefficients (including own-lags) generates spurious cross-market spillovers. When the hybrid approach preserves realistic own-lag persistence via OLS and applies ElasticNet only to cross-terms, most apparent spillover channels vanish. This suggests that many “spillovers” identified in standard regularized VAR models may be artifacts of mis-specified own-dynamics. When a model under-estimates persistence (e.g., shrinking own-lag coefficients from 0.99 to 0.15), it compensates by attributing persistence to cross-market terms—creating false evidence of transmission. The hybrid model corrects this confound, revealing that volatility is predominantly a within-market phenomenon with minimal cross-market linkages in this six-asset system.
4.3 Impulse Response Analysis: Commodities Shock
The network structure identified in Section 4.2 implies specific predictions about how shocks propagate. We now test these predictions using Joint Impulse Response Functions that trace shock dynamics through the estimated network.
Figure 4 shows impulse responses to a simultaneous one-standard-deviation shock to soybean (ZS) and crude oil (CL) volatility. Shaded regions represent 95% bootstrap confidence intervals constructed from 1,000 block bootstrap samples.
@@ 4.3 Impulse Response Analysis: Commodities Shock
The commodity shock generates strong, persistent within-group effects. Crude oil volatility increases by approximately 0.17 (roughly 50% relative to its unconditional mean of 0.34) on impact, decaying slowly to 0.12 by day 20—retaining approximately 71% of the initial shock. This slow decay reflects the high own-lag persistence (around 0.99) preserved by the hybrid model’s OLS HAR step. Soybean volatility shows a smaller initial response of approximately 0.08 (about 39% of its unconditional mean of 0.207), decaying to 0.05 by day 20 (62% of impact remains).
Cross-market spillovers from commodities are negligible. Equity indices, treasury futures, and other asset classes show essentially no response to commodity shocks; confidence intervals include zero throughout the 20-day horizon. This pattern confirms that commodity volatility shocks remain contained within the commodity complex, consistent with the view that commodity markets are “receivers” rather than “transmitters” of volatility in the broader financial system.
4.4 Impulse Response Analysis: Equities Shock
@@
Equity shocks generate strong, highly persistent own-effects but very little cross-market propagation. Both equity indices show responses of approximately 0.095 (S&P 500) and 0.097 (Nasdaq-100) on impact—representing roughly 50-60% relative to their unconditional means and decay slowly over the 20-day horizon, retaining approximately 72% of the initial shock by day 20. This realistic persistence reflects the roughly 0.99 own-lag coefficients preserved by the hybrid model’s OLS HAR step. While we did not observe any statistically significant volatility spillovers, we see a clear expansion of volatility of volatility in the growing width of the crude oil 95% confidence interval response.
4.5 Impulse Response Analysis: Treasuries Shock
The network structure in Section 4.2 shows treasuries primarily receiving rather than transmitting volatility. Figure 6 confirms their peripheral network position, showing impulse responses to a simultaneous shock to 5-year (ZF) and 10-year (ZN) treasury futures volatility.
@@ 4.5 Impulse Response Analysis: Treasuries Shock
Treasury shocks exhibit realistic persistence within the treasury complex but no cross-market transmission making it to the mean response. Both 5-year and 10-year treasury volatility show impact responses near zero, as expected given their low unconditional volatility (mean RV of 0.04 for both). While the absolute magnitudes are small, the relative persistence is consistent with other assets: treasury volatility decays slowly, retaining the characteristic roughly 0.99 persistence from the HAR own-lag structure. Cross-market spillovers are completely absent, but soybeans and crude oil show an increase in volatility of volatility in that the width of the 95% confidence intervals grow over the response horizon.
@@ 5 Conclusion
This paper demonstrates that separating own-volatility dynamics from cross-market spillovers is essential for credible network identification. Using over 20 years of daily realized volatility across six futures markets, we develop a hybrid HARX-ElasticNet framework that estimates own-lag persistence via OLS (preserving the characteristic roughly 0.99 persistence of volatility) before applying ElasticNet regularization to cross-market terms. This two-step procedure reveals an extremely sparse network: only 7 of 90 cross-market spillover coefficients (8%) are nonzero, with equities (ES, NQ) and treasuries (ZF, ZN) exhibiting zero statistically significant cross-market spillovers. The surviving spillovers are concentrated in commodity markets, where crude oil receives small spillovers from equities and treasuries, while soybeans remain largely isolated. We did observe some response in the volatility of volatility as crude oil and soybean JIRF confidence intervals grew over the response horizons when equity or treasuries were shocked.
This extreme sparsity contrasts sharply with standard HARX-ElasticNet, which identifies 40-50% nonzero cross-market coefficients and suggests equities as dominant transmitters. Our methodological contribution is to show that this discrepancy reflects a fundamental issue that must be dealt with to properly identify spillover networks. Uniform regularization applied to all coefficients (including own-lags) shrinks persistence from 0.99 to 0.08-0.18, forcing the model to compensate by attributing persistence to spurious cross-market pathways. The hybrid approach corrects this mis-specification, revealing that volatility is predominantly a within-market phenomenon.
Consistent with Buccheri et al. (2021), the univariate HAR model and our hybrid specification achieve identical out-of-sample RMSE (0.0044), reflecting the hybrid model’s extreme sparsity: with only 8% nonzero cross-market coefficients, it effectively reduces to univariate HAR for forecasting purposes. This demonstrates that cross-market information adds negligible signal for point prediction.
@@
Our results illustrate the broader theme that regularization methods can serve multiple purposes in high-dimensional econometrics. When the goal is forecasting, parsimony typically wins. When the goal is network identification, regularization reveals economically meaningful structure by shrinking noise while preserving signal. The trade-off between these objectives is inherent in the methodology, and practitioners should choose their tools with this in mind.

### v4

@@ 1 Introduction
Third, Joint Impulse Response Functions (JIRFs) illustrate how correlated volatilities can amplify and cross-propagate over time. The JIRF methodology does not identify causal channels; rather, it traces the reduced-form dynamics of joint volatility movements implied by the estimated covariance structure. Despite sparse coefficient estimates, JIRFs reveal substantial co-movement: when equity volatility rises, crude oil volatility tends to rise as well (from 0.04 to 0.09 over 20 days), reflecting the historical correlation between these markets. The accumulating response patterns and widening confidence intervals highlight how correlated shocks can compound through the system, even when direct spillover coefficients are small.
@@ 3.4 HAR Benchmark
3.5 Reduced Form Spillover Network
The advantage of the hybrid HARX-ElasticNet framework is that the estimated coefficient matrix directly reveals the reduced form spillover network structure. The $\gamma$ coefficients indicate which cross-market transmission channels are economically meaningful (those that survive regularization) versus which are noise that gets shrunk to zero.
@@ 4.1 Forecast Performance
4.2 Reduced Form Spillover Network Structure
While HAR excels at point forecasting, it cannot estimate cross-market volatility transmission. The HARX-ElasticNet reveals which spillover pathways are economically meaningful by shrinking noise to zero while preserving significant reduced form cross-market effects. Figure 3 presents the complete estimated coefficient matrix (18 features $\times$ 6 assets = 108 coefficients), where each element represents how lagged volatility in one market (source) at a specific frequency (daily, weekly, monthly) affects current volatility in another market (target).
@@ 4.2 Reduced Form Spillover Network Structure
4.3 Joint Impulse Response Analysis: Commodities Shock
The network structure revealed in Section 4.2 implies specific predictions about how shocks propagate. We now test these predictions using Joint Impulse Response Functions that trace shock dynamics through the estimated network.
Figure 4 shows impulse responses to a simultaneous one-standard-deviation shock to soybean (ZS) and crude oil (CL) volatility. Shaded regions represent 95% bootstrap confidence intervals constructed from 1,000 residual block bootstrap samples.
@@ 4.3 Joint Impulse Response Analysis: Commodities Shock
The commodity shock generates strong, persistent within-group effects that accumulate over the forecast horizon. Crude oil volatility increases by approximately 0.14 on impact and rises to approximately 0.22 by day 20—an accumulation pattern reflecting how correlated volatilities compound through the JIRFs. This rising trajectory, combined with widening confidence intervals, illustrates how joint volatility shocks can amplify over time. Soybean volatility shows a more muted response of approximately 0.06 on impact, remaining relatively stable through day 20, consistent with soybeans’ lower unconditional volatility.
Cross-market co-movements following commodity shocks are small but detectable. Equity indices (ES and NQ) show modest positive responses of approximately 0.02–0.03 that persist throughout the horizon, with confidence intervals remaining above zero. Treasury futures show essentially no response, indicating minimal historical correlation between commodity and treasury volatility. These patterns reflect the reduced-form correlation structure in the data. The asymmetric pattern—positive co-movement with equities but not treasuries—is consistent with the historical tendency for commodity and equity volatility to move together during periods of broad market stress.
4.4 Joint Impulse Response Analysis: Equities Shock
@@
Equity shocks generate strong own-effects that accumulate over the forecast horizon, along with notable co-movement in crude oil. Both equity indices show responses of approximately 0.09 on impact, rising to 0.13–0.14 by day 20. This accumulating pattern, combined with progressively widening confidence intervals, again reflects how correlated volatilities compound through the JIRFs.
The most notable cross-market pattern is the co-movement with crude oil: CL volatility rises from approximately 0.04 on impact to 0.09 by day 20, with confidence intervals that widen substantially and remain above zero throughout the horizon. This pattern reflects the historical correlation between equity and oil volatility—the JIRF does not identify whether equities “cause” oil volatility or whether both respond to common underlying factors. The co-movement is consistent with the observed tendency for equity and energy markets to experience elevated volatility simultaneously during periods of broad market stress. Soybeans and treasuries show minimal response, indicating weaker historical correlation with equity volatility.
4.5 Joint Impulse Response Analysis: Treasuries Shock
The network structure in Section 4.2 shows treasuries primarily transmitting rather than receiving volatility. Figure 6 confirms their peripheral network position, showing impulse responses to a simultaneous shock to 5-year (ZF) and 10-year (ZN) treasury futures volatility.
@@ 4.5 Joint Impulse Response Analysis: Treasuries Shock
Treasury shocks generate the smallest responses of the three shock groups, reflecting the low unconditional volatility of fixed-income markets (mean RV of approximately 0.04 for both ZF and ZN). Own-market responses are modest: 5-year treasury volatility increases by approximately 0.01–0.015, while 10-year treasury volatility shows a response of approximately 0.02, both remaining relatively stable through the forecast horizon. The moderate correlation between ZF and ZN residuals (0.703) generates some within-complex amplification, but the absolute magnitudes remain small.
Despite treasuries’ peripheral network position, we observe small positive co-movement with crude oil and equities. CL volatility rises from approximately 0.02 to 0.04 over the horizon, while ES shows a similar pattern with widening confidence intervals. These treasury-risk-asset correlations, though modest, are consistent with the reduced-form patterns one would expect from interest rate channels: when treasury volatility rises, risk assets tend to experience elevated volatility as well. Soybeans show essentially no response, suggesting that treasury-commodity co-movement is limited to more financially-traded commodities like crude oil.
@@ 5 Conclusion
This paper demonstrates that separating own-volatility dynamics from cross-market spillovers is essential for reduced form volatility spillover network estimation. Using over 20 years of daily realized volatility across six futures markets, we use a hybrid HARX-ElasticNet model that estimates own-lag persistence via OLS (preserving the characteristic roughly 0.99 persistence of volatility) before applying ElasticNet regularization to cross-market terms. This two-step procedure reveals an extremely sparse network: only 7 of 90 cross-market spillover coefficients (8%) are nonzero, with equities (ES, NQ) and treasuries (ZF, ZN) exhibiting zero statistically significant cross-market spillovers in the coefficient estimates. Joint Impulse Response Functions show fairly rich reduced-form dynamics, though. Commodity shocks are associated with small equity volatility increases, equity shocks show notable co-movement with crude oil, and treasury shocks exhibit modest correlations with risk assets. These JIRF patterns demonstrate how sparse point estimates can nonetheless generate meaningful cumulative responses as correlated volatilities compound through the network over time even when many individual coefficients are small.
Consistent with Buccheri et al. (2021), the univariate HAR model and our hybrid specification achieve identical out-of-sample RMSE (0.0044) on average, reflecting the hybrid model’s extreme sparsity: with only 8% nonzero cross-market coefficients, it effectively reduces to univariate HAR for forecasting purposes. This demonstrates that cross-market information adds negligible signal for point prediction.
@@
Our results illustrate the broader theme that regularization methods can serve multiple purposes in high-dimensional econometrics. When the goal is forecasting, parsimony typically wins. When the goal is network identification, regularization can reveal structure. The trade-off between these objectives is inherent in the methodology, and practitioners should choose their tools with this in mind.